# Nonlocal Models in Biology and Life Sciences: Sources, Developments, and Applications

### v1

@@ 4 Emerging Trends, Future Directions, and Outlook
Coupled effects and phenomena, anisotropy, and auxeticity in biosystems: Many biological tissues display features that are dependent on direction, making the appropriate representation of anisotropy in biological tissues essential to understanding their mechanical behaviour. For instance, the human cornea involves capturing its complex structure and mechanical behaviour for various applications, such as understanding biomechanics, predicting responses to surgical procedures, and designing contact lenses. In addition, the modelling of scleral collagen requires the overall tissue architecture, anisotropy, and interactions with other ocular structures. The incorporation of experimental data and imaging investigations into computational nonlocal models can improve their accuracy and applicability mahapatra2006 . Flexoelectricity, piezoelectricity, and auxeticity are coupled effects that can have a major impact on the mechanical and electrical behaviour of materials, particularly in the setting of biosystems singh2022 . Furthermore, the study of piezoelectricity and other coupled effects in biological tissues has gained significant attention due to their potential applications in various biomedical fields melnik1998 ; melnik2000a ; melnik2000b ; melnik2001 ; melnik2003 ; melnik2003a . For instance, smart materials can exhibit both ferroelectric and piezoelectric properties (ferroelectric ceramics). Additionally, multiferroic materials show the coexistence of ferroelectricity and ferromagnetism, and they have been used for different bio applications melnik2000f ; melnik2004 ; matus2004 ; wang2004 ; melnik2004a ; wang2006c ; wang2006d ; wang2007 ; wang2007a ; mahapatra2007a ; wang2007c ; xiao2008 ; wang2009 ; wang2010 ; wang2011 ; he2018 ; wang2016 ; he2020 ; du2020 ; du2020b ; he2020a ; wang2017b ; wang2015 . Shape memory alloys (SMAs) have distinct thermomechanical behaviour, and their hysteresis loops are impacted by the intricate interaction of mechanical and thermal factors. These SMAs have a wide range of biological applications, including thermoelasticity of red blood cell membranes and other biostructures melnik2001a ; melnik2003b ; mahapatra2006a ; wang2006 ; wang2007b ; mahapatra2007 ; wang2008 ; wang2008a ; dhote2012 ; wang2012 ; dhote2013 ; dhote2015 ; dhote2015a ; dhote2016 ; du2021 ; han2022 ; han2022a ; melnik2001 . Nonlocal operations allow each pixel or area in an image to evaluate information from all other places, allowing the model to grasp the broader context fu2019 . This is useful for spotting small structures, deformities, or subtle patterns in biological pictures. These models also help in understanding cellular processes, interactions, and behaviours at the nanoscale conte2023 . Fractional calculus presents the notion of fractional order derivatives, which enables the modelling of processes with memory effects, long-range interactions, and nonlocal behaviour. This allows us to have a more accurate representation of biological and nanoscale processes with complex and nonlocal dynamics metzler1994 .
@@
Nonlocal models for advancing the frontiers of medicine, nanoscience and nanotechnology: Nonlocal models provide a more accurate and comprehensive understanding of the behaviour of materials at the nanoscale. Nanocomposites are composed materials, a combination of nanoparticles and matrix material, which enhances their functionalities and opens up new possibilities in fields such as medicine, diagnostics, and biotechnology fallahpour2022 . Self-assembled bio-nanostructures, like RNA-nanostructures, have a large number of applications across various fields mahapatra2012 ; bahrami2010 ; prabhakar2010 ; patil2009 ; melnik2007 . These can be designed to encapsulate and deliver therapeutic agents like drugs or gene therapies to specific target cells or even in the development of RNA-based vaccines paliy2010 ; paliy2009 ; badu2014 ; badu2020 ; badu2020a ; badu2020b ; badu2020c ; badu2020d ; badu2015 ; badu2015a ; badu2015b ; melnik2002b . Different biotechnologies based on RNA and DNA collectively contribute to advancements in medicine, basic research, and various industries yang2007 ; yang2007b . The nano-template viruses can be modified to carry therapeutic agents to specific cells or tissues, enhancing drug delivery precision and reducing off-target effects. In addition, carbon-based nanomaterials exhibit remarkable properties and have been used in emerging technology with potential applications in medical imaging and other fields sinha2007 ; sinha2007a ; mahapatra2008 ; roy2009 ; mahapatra2011 ; yavarian2023 . Biosensors detect specific analytes and convert this information into a measurable signal singh2022a . Nevertheless, environmentally-friendly technologies (e.g., biomedical and chemical sensors) for life sciences play crucial roles in various applications, providing valuable information for monitoring, diagnosis, and research singh2022a ; fallahpour2021 . As materials are scaled down to the nanoscale, size-dependent effects and phenomena become more evident, e.g., flexoelectric effects are becoming more pronounced at the nanoscale. For instance, the nanoscale structures have many applications in biomedical and bioengineering, such as nanowires, nanorods, and nanowire superlattices lew2004 ; lassen2004 ; willatzen2004 ; galeriu2004 ; lew2004 ; radulovic2006 ; lassen2006a ; patil2009a ; patil2009b ; alvaro2013 ; prabhakar2013 and quantum dots melnik2004 ; melnik2003f ; melnik2005 ; lew2005 ; prabhakar2012 ; prabhakar2012a ; prabhakar2013a . Furthermore, it is important to include the strain and the coupled effects together lassen2006 ; willatzen2006 ; lassen2005 ; lassen2005a . The utilization of 2D materials such as graphene with pronounced coupled effects in biomedical engineering, e.g., graphene nanoribbons prabhakar2013a ; prabhakar2015 ; prabhakar2016 ; prabhakar2017 ; prabhakar2019 ; leon2022 , and other graphene-based nanostructures prabhakar2014 ; prabhakar2015 ; prabhakar2015a ; prabhakar2016 ; prabhakar2013b ; prabhakar2022 .
@@
Nonlocal neural models: Neurons in the brain communicate and interact with each other through synapses, which can involve long-range connections lovinger2008 ; seguin2023 ; papo2024 . Nonlocal neural models have been employed to describe neurons’ collective activity and synchronization in neural networks. These models are gaining attention to represent neural networks’ long-range interactions and memory effects heffernan2017 . They can provide insights into brain function, learning, and cognition, e.g., how the information is processed and transmitted between neurons in neural networks, and they can benefit from considering nonlocal interactions.
@@
Fractional model: Ecological systems often span multiple scales, from individual organisms to entire ecosystems. Fractional reaction-diffusion models can account for nonlocal interactions and dispersal across different spatial scales, making them valuable tools for studying population dynamics and species interactions in diverse ecological settings owolabi2023 . In addition, the study of anomalous diffusion, which departs from classical diffusion behaviour, is becoming more prevalent in nonlocal models. Fractional diffusion and subdiffusion models are being used to describe transport processes in complex biological media with irregular geometries or barriers obembe2020 . For instance, the drug delivery within the tumor often exhibits anomalous transport behaviour.
@@
Nonlocal models in ecology: Ecological systems are defined by complex interactions between species at various spatial scales lang2013 ; berrio2021 . Nonlocal models are used to investigate species dispersal, habitat connectivity, and the effect of environmental changes on population dynamics. Furthermore, depending on the nature of the interactions, several types of nonlocal interaction are considered in the reaction terms, as well as self- or cross-diffusion jdmurray ; volpert2009 ; pal2018bmb . Nonlocal dispersal is an essential aspect of ecological processes, and understanding its implications is vital for predicting species distributions, ecosystem dynamics, and responses to environmental changes sherratt2016 . As ecological research advances, the study of nonlocal dispersal in biological systems will likely remain a critical investigation area.
@@
Treating diseases with innovative procedures: It involves exploring cutting-edge techniques, technologies, and approaches to improve patient outcomes. Some examples of innovative procedures in disease treatment include cryosurgery - cell destruction, thermal ablation, and laser ablation singh2020 ; zhang2008 ; wang2006a ; wang2006b . Phase-lag models are mathematical frameworks used to describe the heat transfer processes in biological tissues. These models are useful for understanding thermal processes in live beings and are widely used in thermal therapy, hyperthermia therapies, and thermal imaging singh2021 ; singh2022b ; singh2020d ; singh2020e . Cardiac ablation, in particular, is a medical technique used to treat irregular heart rhythms by selectively eliminating or isolating portions of the heart tissue that create or conduct irregular electrical impulses singh2019 ; singh2019a ; singh2019b ; singh2020f .
@@
Brain dynamics and modelling in neuroscience: It is an interdisciplinary proposal to discover how the brain creates cognition and behaviour. Neuroimaging methods, such as fMRI and EEG, offer data for the development of data-driven models capable of capturing the spatiotemporal dynamics of brain activity and connection. These models are critical for creating brain-machine interfaces, which transform cerebral activity into commands for external devices, allowing persons with motor disabilities to communicate and operate technology dayan2005 . Nonlocal models provide a framework for representing the long-distance connections that contribute to information integration and processing, as neural interactions frequently extend beyond immediate neighbours. Furthermore, brain dynamic models aid in the understanding of aberrant brain dynamics associated with illnesses such as Alzheimer’s disease and Parkinson’s disease pal2022sr ; shaheen2022 ; shaheen2023 .
@@
Nonlocal models on network: Nonlocal interactions on networks have broader applications in systems biology, including protein-protein interaction networks, cell signalling networks, and ecological food webs autry2018 ; sevimoglu2014 ; lee2022 . Understanding and incorporating nonlocal interactions on networks are essential for capturing biological systems’ emergent properties and collective behaviours. As research in network biology progresses, the significance of nonlocal interactions in shaping the dynamics and functions of biological networks will continue to be a topic of interest and investigation.
@@
Nonlocal models in personalized medicine: Nonlocal models are being investigated in medicine to explain better individual variability and tailored responses to treatments falco2021 . Individual-specific dynamics in diseases are being characterized using fractional calculus, and individualized therapeutic actions are being predicted.
@@
Nonlocal models in epidemiology: Nonlocal epidemiological models are valuable in understanding disease dynamics in real-world scenarios, where the assumption of local interactions is often inadequate ruan2007 . They have applications in public health planning, disease control strategies, and outbreak predictions. As our understanding of infectious diseases and computational methods continue to advance, nonlocal epidemiological models will play an increasingly crucial role in addressing emerging challenges in biology and epidemiology.
@@
Nonlocal interactions in social networks and biosocial dynamics: Complex networks provide a powerful lens for understanding the structure and behaviour of interconnected systems in diverse domains melnik2008 ; dankulov2019 . For instance, social networks exhibit nonlocal interactions which extend connections among individuals. The structure of social networks is characterized by ties that go beyond direct acquaintances and influence the spread of information, behaviours, and opinions tadic2017 ; newman2018 ; dankulov2015 ; andjelkovic2016 . On the other hand, biosocial dynamics incorporate the spread of health behaviours, such as smoking cessation, exercise adoption, and vaccination uptake christakis2007 . The connectivity patterns within populations influence the dynamics of disease transmission, and network models assist in designing effective control strategies tadic2021 ; tadic2020 ; mitrovic2023 ; mitrovic2022 . Human connectomes are highly connected hubs involving complex interactions between regions and their relevance for different processes tadic2019 ; andjelkovic2020 . Furthermore, nonlocal interactions contribute to the transmission of emotional experiences, influencing mental health and well-being kramer2014 .
@@
Conservation laws in biological transport: The principles of conservation laws, such as mass or energy conservation, are fundamental in understanding various biological transport processes sterner2011 . For instance, they are applied in modelling blood flow in vessels, nutrient transport in plants, and fluid dynamics in biological systems sharma2018 ; el2022 . By employing nonlocal problems and conservation laws in mathematical modelling, researchers can better understand complex biological phenomena and develop more accurate and realistic representations of biological processes involving long-range interactions and mass conservation.
@@
Nonlocal models in integrating multiscale modelling and AI techniques: Nonlocal models play a crucial role in integrating multiscale modelling and AI techniques, providing a framework that captures interactions and influences beyond immediate neighbours or specific scales. For instance, AI techniques in neural networks can learn and represent long-range interactions, which allows for more accurate predictions and simulations across different scales. In addition, AI techniques can assist in optimizing parameters across different scales based on nonlocal information, improving the overall fidelity of the multiscale model. There are mathematical and computational challenges in the modelling of complex systems that involve multiphysics and multiscale processes du2020 . But, overall, nonlocal models serve as a bridge in integrating multiscale modelling and AI techniques by providing a comprehensive framework for capturing new insights into disease mechanisms on different scales, which helps in identifying new targets and treatment strategies for the benefit of human health alber2019 .
@@
Active matter and nonlocal models: Active matter refers to a class of materials or systems that can convert energy into mechanical work to generate self-driven motion. These systems are distinguished by their individual components’ capacity to drive themselves actively, showing aggregate behaviours that result from interactions between these active entities. Active matter is a large and multidisciplinary area that includes biological creatures, synthetic microswimmers, and even certain types of materials ghosh2020 ; gompper2020 . Nonlocality is essential for capturing the long-range correlations and collective behaviours that appear often in active matter systems. For example, the Vicsek model depicts the collective motion of self-propelled particles aligning with their neighbours vicsek1995 . It considers the average alignment of particles in a certain neighbourhood to add nonlocal interactions. Several theoretical frameworks and mathematical models, generally taken from statistical mechanics, fluid dynamics, and soft matter physics, have been constructed to characterize active matter and integrate nonlocal interactions. This method is critical for researching phenomena like flocking, swarming, and pattern development.
@@
Nonlocal models in bio-inspired technologies: Nonlocal models help in capturing the nonlocal interactions and behaviours observed in biological systems, which allows for the development of more efficient, adaptive, and biomimetic technologies. These models are employed in simulating the adaptive properties of biological materials and structures, e.g., self-healing, self-repair, and responsiveness to environmental changes. Furthermore, these models allow for the simulation of how bio-inspired technology often involves the development of sensors for enhancing sensory systems in areas including environmental monitoring and surveillance. These emerging developments emphasize the growing importance of nonlocal mathematical models in solving complicated biological challenges and exploring new mathematical biology frontiers. In addition, electron transfer processes are vital in many chemical reactions and biological systems. Nonlocal models, such as the Marcus theory, consider the nonlocal electron transfer interactions between donor and acceptor molecules. These models account for the influence of solvent polarization and nuclear motions on electron transfer rates, leading to a more accurate description of electron transfer phenomena. Nonlocal models are likely to play a critical role in enhancing our understanding of varied biological systems as interdisciplinary research expands.
@@
Complex adaptive systems: There is a growing interest in complex adaptive systems, and a new framework is being created to combine classical and quantum systems using “emergent” quantum theory in complex adaptive systems paperin2011 ; hubsch2023 . Some hidden variables in classical nonlocal models impact the outcomes of measurements and can reveal correlations that local, realistic theories cannot explain. Quantum mechanics, on the other hand, introduces nonlocality through phenomena such as entanglement, in which particles become correlated. Nonlocality is a fundamental property of quantum physics, although it may also be investigated in some classical systems. Interestingly, the complex adaptive system is directly relevant to Lotka-Volterra systems, and researchers argue that the emergent quantum theory of Lotka-Volterra dynamics is described by the emergent stationary Schrodinger equation.
@@
Coarse-grained models and computations with equation-free models: Coarse-grained models are simplified representations of complicated systems that average or combine small components to capture the essential aspects. These models involve constructing a hierarchy of models that explain the system at various degrees of detail. On the other hand, equation-free models are computational tools that allow for the study and simulation of complex systems without explicit knowledge of governing equations. These strategies are especially effective when the system dynamics are unclear or computationally costly. Equation-free models are used in a wide range of disciplines, including neurology, climate science, and materials research. Coarse-grained models and equation-free models are both effective tools for modelling and evaluating complex systems in situations when standard methodologies are unfeasible, according to gear2003 . They enable researchers to get insights into the emergent behaviour and dynamics of multiscale systems without having to study the microscopic processes in depth.
@@
Hyperbolic and thermal relaxation models, other sources of nonlocality: Hyperbolic thermoelasticity is the coupled of heat conduction and elastic wave equations and describes the evolution of temperature and displacement fields in the material. It has many applications in various fields, including biology, the design of thermal devices, and the understanding of heat transfer mechanisms at the microscopic level strunin2000 ; strunin2001 ; hillen2002 ; melnik2002a ; melnik2005 . In addition, the heat conduction equation with appropriate relaxation terms captures the time-dependent behaviour of the material during thermal ablation singh2019d , which is typically used to destroy or damage tissues for the treatment of tumors. Furthermore, nonlocal PDEs play a significant role in modelling various phenomena, e.g., swarming behaviour, nonlinear diffusion, models in fluid dynamics and fluid-structure interaction for biological and life science problems liang2019 ; kamath2006 . Nonlinear nonlocal models also account for competing effects (e.g., attractive and repulsive forces). Biological tissues, such as tendons, ligaments, and bone, have intricate microstructures that can lead to non-uniform stress and strain distributions. Nonlocal plasticity models are motivated by the need to capture size effects, strain localization, and the influence of remote material points on deformation patterns in these tissues, enhancing our understanding of tissue mechanics and guiding advancements in bioengineering nilsson1998 ; shizawa1999 ; lion2000 . In addition, phase-field models in materials are used to study the evolution of phase boundaries, such as solidification, melting, and phase separation provatas2011 ; dhote2014 ; dhote2014a ; dhote2014b ; wang2017a ; wang2018a . The nonlocal effects in the phase-field models are especially relevant in materials where the interface width is not negligible compared to the system size. Contact mechanics also considers the nonlocal model, which deals with the deformation and interaction of solid surfaces in contact. Nevertheless, the spatial organization and arrangement of certain features or structures happens within a semiconductor material scholl2001 . This phenomenon is significant in semiconductor technology as it directly affects the operation and performance of electronic devices, including those critical in healthcare bioengineering and biomedicine.

### v2

@@ 4 Emerging Trends, Future Directions, and Outlook
Coupled effects and phenomena, anisotropy, and auxeticity in biosystems: Many biological tissues display features dependent on direction, making the appropriate representation of anisotropy in biological tissues essential to understanding their mechanical behaviour. For instance, the human cornea involves capturing its complex structure and mechanical behaviour for various applications, such as understanding biomechanics, predicting responses to surgical procedures, and designing contact lenses. In addition, the modelling of scleral collagen requires the overall tissue architecture, anisotropy, and interactions with other ocular structures. Incorporating experimental data and imaging investigations into computational nonlocal models can improve their accuracy and applicability [517]. Flexoelectricity, piezoelectricity, and auxeticity are coupled effects that can have a major impact on the mechanical and electrical behaviour of materials, particularly in the setting of biosystems [518]. Furthermore, the study of piezoelectricity and other coupled effects in biological tissues has gained significant attention due to their potential applications in various biomedical fields [519, 520, 521, 522, 523, 524]. For instance, smart materials can exhibit ferroelectric and piezoelectric properties (ferroelectric ceramics). Additionally, multiferroic materials show the coexistence of ferroelectricity and ferromagnetism, and they have been used for different bio applications [525, 526, 527, 528, 529, 530, 531, 532, 533, 534, 535, 536, 537, 538, 539, 540, 541, 542, 543, 544, 545, 546, 369]. Shape memory alloys (SMAs) have distinct thermomechanical behaviour, and the intricate interaction of mechanical and thermal factors impacts their hysteresis loops. These SMAs have a wide range of biological applications, including thermoelasticity of red blood cell membranes and other biostructures [547, 548, 549, 550, 551, 552, 553, 554, 555, 556, 557, 558, 559, 560, 561, 562, 563, 522]. Nonlocal operations allow each pixel or area in an image to evaluate information from all other places, allowing the model to grasp the broader context [564]. This is useful for spotting small structures, deformities, or subtle patterns in biological pictures. These models also help to understand cellular processes, interactions, and behaviours at the nanoscale [565]. Fractional calculus presents the notion of fractional order derivatives, which enables the modelling of processes with memory effects, long-range interactions, and nonlocal behaviour. This allows us to more accurately represent biological and nanoscale processes with complex and nonlocal dynamics [566].
@@
Nonlocal models for advancing the frontiers of medicine, nanoscience and nanotechnology: Nonlocal models provide a more accurate and comprehensive understanding of the behaviour of materials at the nanoscale. Nanocomposites are composed materials, a combination of nanoparticles and matrix material, which enhances their functionalities and opens up new possibilities in fields such as medicine, diagnostics, and biotechnology [567]. Self-assembled bio-nanostructures, like RNA-nanostructures, have a large number of applications across various fields [568, 569, 570, 571, 572]. These can be designed to encapsulate and deliver therapeutic agents like drugs or gene therapies to specific target cells or even in the development of RNA-based vaccines [494, 493, 495, 496, 497, 498, 573, 574, 575, 576, 577, 578]. Different biotechnologies based on RNA and DNA collectively contribute to advancements in medicine, basic research, and various industries [111, 579]. The nano-template viruses can be modified to carry therapeutic agents to specific cells or tissues, enhancing drug delivery precision and reducing off-target effects. In addition, carbon-based nanomaterials exhibit remarkable properties and have been used in emerging technology with potential applications in medical imaging and other fields [580, 581, 582, 583, 584, 585]. Biosensors detect specific analytes and convert this information into a measurable signal [586]. Nevertheless, environmentally-friendly technologies (e.g., biomedical and chemical sensors) for life sciences play crucial roles in various applications, providing valuable information for monitoring, diagnosis, and research [586, 587]. As materials are scaled down to the nanoscale, size-dependent effects and phenomena become more evident, e.g., flexoelectric effects are becoming more pronounced at the nanoscale. For instance, the nanoscale structures have many applications in biomedical and bioengineering, such as nanowires, nanorods, and nanowire superlattices [588, 589, 590, 591, 588, 592, 593, 594, 595, 596, 597] and quantum dots [526, 598, 599, 600, 601, 602, 603]. Furthermore, including the strain and the coupled effects together [604, 605, 606, 607]. The utilization of 2D materials such as graphene with pronounced coupled effects in biomedical engineering, e.g., graphene nanoribbons [603, 608, 609, 610, 611, 612], and other graphene-based nanostructures [613, 608, 614, 609, 615, 616].
@@
Nonlocal neural models: Neurons in the brain communicate and interact through synapses, which can involve long-range connections [617, 618, 619]. Nonlocal neural models have been employed to describe neurons’ collective activity and synchronization in neural networks. These models are gaining attention to represent neural networks’ long-range interactions and memory effects [620]. They can provide insights into brain function, learning, and cognition, e.g., how the information is processed and transmitted between neurons in neural networks, and they can benefit from considering nonlocal interactions.
@@
Fractional model: Ecological systems often span multiple scales, from individual organisms to entire ecosystems. Fractional reaction-diffusion models can account for nonlocal interactions and dispersal across different spatial scales, making them valuable tools for studying population dynamics and species interactions in diverse ecological settings [621]. In addition, the study of anomalous diffusion, which departs from classical diffusion behaviour, is becoming more prevalent in nonlocal models. Fractional diffusion and subdiffusion models are being used to describe transport processes in complex biological media with irregular geometries or barriers [622]. For instance, the drug delivery within the tumor often exhibits anomalous transport behaviour.
@@
Nonlocal models in ecology: Ecological systems are defined by complex interactions between species at various spatial scales [623, 624]. Nonlocal models are used to investigate species dispersal, habitat connectivity, and the effect of environmental changes on population dynamics. Furthermore, depending on the nature of the interactions, several types of nonlocal interaction are considered in the reaction terms, as well as self- or cross-diffusion [4, 72, 88, 625, 626, 627, 628, 629]. Nonlocal dispersal is an essential aspect of ecological processes, and understanding its implications is vital for predicting species distributions, ecosystem dynamics, and responses to environmental changes [54]. As ecological research advances, the study of nonlocal dispersal in biological systems will likely remain a critical investigation area.
@@
Treating diseases with innovative procedures: It involves exploring cutting-edge techniques, technologies, and approaches to improve patient outcomes. Some examples of innovative processes in disease treatment include cryosurgery - cell destruction, thermal ablation, and laser ablation [118, 630, 631, 632]. Phase-lag models are mathematical frameworks which describe the heat transfer processes in biological tissues. These models are useful for understanding thermal processes in live beings and are widely used in thermal therapy, hyperthermia therapies, and thermal imaging [633, 634, 635, 636]. Cardiac ablation, in particular, is a medical technique used to treat irregular heart rhythms by selectively eliminating or isolating portions of the heart tissue that create or conduct irregular electrical impulses [637, 638, 639, 640].
@@
Brain dynamics and modelling in neuroscience: It is an interdisciplinary proposal to discover how the brain creates cognition and behaviour. Neuroimaging methods, such as fMRI and EEG, offer data for developing data-driven models to capture the spatio-temporal dynamics of brain activity and connection. These models are critical for creating brain-machine interfaces, which transform cerebral activity into commands for external devices, allowing persons with motor disabilities to communicate and operate technology [641]. Nonlocal models provide a framework for representing the long-distance connections that contribute to information integration and processing, as neural interactions frequently extend beyond immediate neighbours. Furthermore, brain dynamic models aid in the understanding of aberrant brain dynamics associated with illnesses such as Alzheimer’s disease and Parkinson’s disease [46, 642, 643].
@@
Nonlocal models on network: Nonlocal interactions on networks have broader applications in systems biology, including protein-protein interaction networks, cell signalling networks, and ecological food webs [175, 644, 645]. Understanding and incorporating nonlocal interactions on networks are essential for capturing biological systems’ emergent properties and collective behaviours. As research in network biology progresses, the significance of nonlocal interactions in shaping the dynamics and functions of biological networks will continue to be a topic of interest and investigation.
@@
Nonlocal models in personalized medicine: Nonlocal models are being investigated in medicine to explain better individual variability and tailored responses to treatments [646]. Individual-specific dynamics in diseases are being characterized using fractional calculus, and individualized therapeutic actions are being predicted.
@@
Nonlocal models in epidemiology: Nonlocal epidemiological models are valuable in understanding disease dynamics in real-world scenarios, where the assumption of local interactions is often inadequate [134]. They have applications in public health planning, disease control strategies, and outbreak predictions. As our understanding of infectious diseases and computational methods continue to advance, nonlocal epidemiological models will play an increasingly crucial role in addressing emerging challenges in biology and epidemiology.
@@
Nonlocal interactions in social networks and biosocial dynamics: Complex networks provide a powerful lens for understanding the structure and behaviour of interconnected systems in diverse domains [647, 492]. For instance, social networks exhibit nonlocal interactions which extend connections among individuals. The structure of social networks is characterized by ties that go beyond direct acquaintances and influence the spread of information, behaviours, and opinions [230, 648, 649, 650]. On the other hand, biosocial dynamics incorporate the spread of health behaviours, such as smoking cessation, exercise adoption, and vaccination uptake [651]. The connectivity patterns within populations influence the dynamics of disease transmission, and network models assist in designing effective control strategies [652, 653, 186, 654]. Human connectomes are highly connected hubs involving complex interactions between regions and their relevance for different processes [655, 656]. Furthermore, nonlocal interactions contribute to the transmission of emotional experiences, influencing mental health and well-being [657].
@@
Conservation laws in biological transport: The principles of conservation laws, such as mass or energy conservation, are fundamental in understanding various biological transport processes [397]. For instance, they are applied in modelling blood flow in vessels, nutrient transport in plants, and fluid dynamics in biological systems [658, 659]. By employing nonlocal problems and conservation laws in mathematical modelling, researchers can better understand complex biological phenomena and develop more accurate and realistic representations of biological processes involving long-range interactions and mass conservation.
@@
Nonlocal models in integrating multiscale modelling and AI techniques: Nonlocal models play a crucial role in integrating multiscale modelling and AI techniques, providing a framework that captures interactions and influences beyond immediate neighbours or specific scales. For instance, AI techniques in neural networks can learn and represent long-range interactions, which allows for more accurate predictions and simulations across different scales. In addition, AI techniques can assist in optimizing parameters across different scales based on nonlocal information, improving the overall fidelity of the multiscale model [660]. There are mathematical and computational challenges in the modelling of complex systems that involve multiphysics and multiscale processes [543]. But, overall, nonlocal models serve as a bridge in integrating multiscale modelling and AI techniques by providing a comprehensive framework for capturing new insights into disease mechanisms on different scales, which helps in identifying new targets and treatment strategies for the benefit of human health [661].
@@
Active matter and nonlocal models: Active matter refers to a class of materials or systems that can convert energy into mechanical work to generate self-driven motion. These systems are distinguished by their individual components’ capacity to drive themselves actively, showing aggregate behaviours that result from interactions between these active entities. Active matter is a large and multidisciplinary area that includes biological creatures, synthetic microswimmers, and even certain types of materials [662, 663]. Nonlocality is essential for capturing the long-range correlations and collective behaviours that appear often in active matter systems. For example, the Vicsek model depicts the collective motion of self-propelled particles aligning with their neighbours [664]. It considers the average alignment of particles in a certain neighbourhood to add nonlocal interactions. Several theoretical frameworks and mathematical models, generally taken from statistical mechanics, fluid dynamics, and soft matter physics, have been constructed to characterize active matter and integrate nonlocal interactions. This method is critical for researching phenomena like flocking, swarming, and pattern development.
@@
Nonlocal models in bio-inspired technologies: Nonlocal models help capture the nonlocal interactions and behaviours observed in biological systems, allowing for the development of more efficient, adaptive, and biomimetic technologies. These models are employed in simulating the adaptive properties of biological materials and structures, e.g., self-healing, self-repair, and responsiveness to environmental changes. Furthermore, these models allow for the simulation of how bio-inspired technology often involves the development of sensors for enhancing sensory systems in areas including environmental monitoring and surveillance. These emerging developments emphasize the growing importance of nonlocal mathematical models in solving complicated biological challenges and exploring new mathematical biology frontiers. In addition, electron transfer processes are vital in many chemical reactions and biological systems. Nonlocal models, such as the Marcus theory, consider the nonlocal electron transfer interactions between donor and acceptor molecules. These models account for the influence of solvent polarization and nuclear motions on electron transfer rates, leading to a more accurate description of electron transfer phenomena. Nonlocal models are likely to play a critical role in enhancing our understanding of varied biological systems as interdisciplinary research expands.
@@
Complex adaptive systems: There is a growing interest in complex adaptive systems, and a new framework is being created to combine classical and quantum systems using “emergent” quantum theory in complex adaptive systems [665, 666]. Some hidden variables in classical nonlocal models impact the outcomes of measurements and can reveal correlations that local, realistic theories cannot explain. Quantum mechanics, on the other hand, introduces nonlocality through phenomena such as entanglement, in which particles become correlated. Nonlocality is a fundamental property of quantum physics, although it may also be investigated in some classical systems. Interestingly, the complex adaptive system is directly relevant to Lotka-Volterra systems, and researchers argue that the emergent stationary Schrodinger equation describes the emergent quantum theory of Lotka-Volterra dynamics.
@@
Coarse-grained models and computations with equation-free models: Coarse-grained models are simplified representations of complicated systems that average or combine small components to capture the essential aspects. These models involve constructing a hierarchy of models that explain the system at various degrees of detail. On the other hand, equation-free models are computational tools that allow for the study and simulation of complex systems without explicit knowledge of governing equations. These strategies are especially effective when the system dynamics are unclear or computationally costly. Equation-free models are used in a wide range of disciplines, including neurology, climate science, and materials research. Coarse-grained models and equation-free models are both effective tools for modelling and evaluating complex systems in situations when standard methodologies are unfeasible, according to [667]. They enable researchers to get insights into the emergent behaviour and dynamics of multiscale systems without having to study the microscopic processes in depth.
@@
Hyperbolic and thermal relaxation models, other sources of nonlocality: Hyperbolic thermoelasticity is the coupled of heat conduction and elastic wave equations and describes the evolution of temperature and displacement fields in the material. It has many applications in various fields, including biology, the design of thermal devices, and the understanding of heat transfer mechanisms at the microscopic level [668, 669, 670, 671, 599]. In addition, the heat conduction equation with appropriate relaxation terms captures the time-dependent behaviour of the material during thermal ablation [672], which is typically used to destroy or damage tissues for the treatment of tumors. Furthermore, nonlocal PDEs play a significant role in modelling various phenomena, e.g., swarming behaviour, nonlinear diffusion, models in fluid dynamics and fluid-structure interaction for biological and life science problems [673, 279]. Nonlinear nonlocal models also account for competing effects (e.g., attractive and repulsive forces). Biological tissues, such as tendons, ligaments, and bone, have intricate microstructures that can lead to non-uniform stress and strain distributions. Nonlocal plasticity models are motivated by the need to capture size effects, strain localization, and the influence of remote material points on deformation patterns in these tissues, enhancing our understanding of tissue mechanics and guiding advancements in bioengineering [674, 675, 676]. In addition, phase-field models in materials are used to study the evolution of phase boundaries, such as solidification, melting, and phase separation [677, 678, 679, 680, 681, 682]. The nonlocal effects in the phase-field models are especially relevant in materials where the interface width is not negligible compared to the system size. Contact mechanics also considers the nonlocal model, which deals with the deformation and interaction of solid surfaces in contact. Nevertheless, the spatial organization and arrangement of certain features or structures happens within a semiconductor material [683]. This phenomenon is significant in semiconductor technology as it directly affects the operation and performance of electronic devices, including those critical in healthcare bioengineering and biomedicine.